Two particles each of mass m and charge q are separated by r1 and the system is left free to move at t=0 At t=t both the particles are found to be separated by r2. The speed of each particle is?

Grade:12th pass

1 Answers

Arun
4 years ago
 
repulsive electrostatic force:  attractive gravitational force:
F1=kQ^2/r^2   F2=Gm^2/r^2
for equilibrium :F1=F2
kQ2/r2=Gm2/r2
=> Q/m = (6.6*10^-11/9*10^9)1/2
=> Q/m=0.8*10-10
Hence, ORDER OF Q/m=10-10